# Policy Mirror Descent with Lookahead

### v1

@@ 8 Related Work
Policy Mirror Descent and Policy Gradient Methods. Motivated by their empirical success (Schulman et al., 2017a, b), the analysis of PG methods has recently attracted a lot of attention (Agarwal et al., 2021; Bhandari and Russo, 2021; Khodadadian et al., 2022; Bhandari and Russo, 2024; Xiao, 2022; Lan, 2023; Johnson et al., 2023).
@@
Among these works, a line of research focused on the analysis of PMD (Xiao, 2022; Lan, 2023; Li and Lan, 2023; Li et al., 2023; Johnson et al., 2023; Alfano et al., 2023) as a flexible algorithmic framework, and its particular instances such as the celebrated natural policy gradient (Khodadadian et al., 2022; Yuan et al., 2023). In particular, for solving unregularized MDPs (which is our main focus), Xiao (2022) established linear convergence of PMD in the tabular setting with a rate depending on the instance dependent distribution mismatch coefficients, and extended their results to the inexact setting using a generative model. Similar results were established for the particular case of the natural policy gradient algorithm (Bhandari and Russo, 2021; Khodadadian et al., 2022). More recently, Johnson et al. (2023) showed a dimension-free rate for PMD with exact policy evaluation using adaptive stepsizes, closing the gap with PI. Notably, they further show that this rate is optimal and their analysis inspired from PI circumvents the use of the performance difference lemma which was prevalent in prior work. While our analysis shares some similarity with this recent work in the tabular setting as we use similar adaptive stepsizes, we improve over these results using lookahead and we extend our analysis beyond the tabular setting by using linear function approximation. To the best of our knowledge, the use of multi-step greedy policy improvement in PMD and its analysis are novel in the literature.
@@
Policy Iteration with Multiple-Step Policy Improvement. As mentioned, multi-step greedy policies have been succesfully used in applications such as the game of Go and have been studied in a body of works (Efroni et al., 2018a, b; Winnicki and Srikant, 2023; Efroni et al., 2020; Tomar et al., 2020). In particular, multi-step greedy policy improvement has been studied in conjunction with PI in Efroni et al. (2018a); Winnicki and Srikant (2023). (Efroni et al., 2018a) introduced $h$-PI which incorporates $h$-lookahead to PI and generalized existing analyses on PI with 1-step greedy policy improvement to $h$-step greedy policies. Our present work builds on these developments together with the connection between PMD and PI. Later, Efroni et al. (2018b) proposed softly and hardly updated multiple-step greedy policies for online and approximate PI. Inspired by this line of works, the class of $h$-PMD algorithms we introduce is novel to the best of our knowledge. Winnicki and Srikant (2023) showed that that a first-visit version of a PI scheme using a single sample path for policy evaluation converges to the optimal policy provided that the policy improvement step uses lookahead instead of a $1$-step greedy policy. Unlike our present work, it is assumed in the latter work that we have access to a lookahead greedy policy oracle.
@@
Lookahead-Based Policy Improvement with Tree Search Lookahead policy improvement is usually implemented via tree search methods (Browne et al., 2012; Shah et al., 2020; Schrittwieser et al., 2019; Kocsis and Szepesvári, 2006b; Hallak et al., 2023; Efroni et al., 2019). Recently, Dalal et al. (2022) proposed a method combining policy gradient and tree search to reduce the variance of stochastic policy gradients. The method relies on a softmax policy parametrization incorporating a tree expansion. Our general $h$-PMD framework which incorporates a different lookahead policy improvement step also brings together PG and tree search methods. Morimura et al. (2022) designed a method with a mixture policy of PG and MCTS for non-Markov Decision Processes. Grill et al. (2020) showed that AlphaZero as well as several other MCTS algorithms compute approximate solutions to a family of regularized policy optimization problems which bear similarities with our $h$-PMD algorithm update rule for $h=1$ (see sec. 3.2 therein).
@@ 9 Conclusion
In this work, we introduced a novel class of PMD algorithms with lookahead inspired by the success of lookahead policies in practice as well as their analysis in PI with multi-step greedy improvements. We have shown an improved $\gamma^{h}$-linear convergence rate depending on the lookahead depth in the exact setting. We proposed an inexact version of our algorithm estimating the lookahead value function via planning and Monte Carlo sampling and we derived an improved sample complexity. We further extended our results to scale to large state spaces via linear function approximation with a performance bound independent on the state space size. Our paper offers several interesting directions for future research. A possible avenue for future work is to investigate the use of more general function approximators such as neural networks to approximate the lookahead value functions and scale to large state-action spaces. Recent work for PMD along these lines (Alfano et al., 2023) might be a good starting point. Enhancing our algorithm with exploration mechanisms (see e.g. the recent work O’Donoghue (2023)) is an important future direction. Our PMD update rule might offer some advantage in this regard as was recently observed in the literature (Li and Lan, 2023). Constructing and studying more efficient estimators for the lookahead action values using for example Monte Carlo Tree Search is also certainly a promising avenue for future work. Designing adaptive lookahead strategies to select the tree search horizon (Rosenberg et al., 2023) is yet another opportunity for further research. Our algorithm brings together two popular and successful families of RL algorithms: PG methods and tree search methods. Looking forward, we hope our work further stimulates research efforts to investigate practical and efficient methods enjoying the benefits of both classes of methods in solving practical RL problems.
@@ Appendix A Additional Details and Proofs for Section 4: Exact $h$-PMD
$=({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\sum_{i=1}^{n}\gamma^{i+1}c_{k}\mathbf{e}+\gamma c_{k}\mathbf{e}$
$=({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\sum_{i=1}^{n+1}\gamma^{i}c_{k}\mathbf{e}\,,$ (14)
@@ Appendix B Additional Details and Proofs for Section 5: Inexact $h$-PMD
B.1 Proof of Theorem 5.3: Inexact $h$-PMD
@@
$\lVert\hat{V}^{{\pi_{k}}}_{h}(\pi)-{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\rVert_{\infty}=\max_{s\in\mathcal{S}}|\langle\hat{Q}_{h}^{\pi_{k}}(s,\cdot)-Q_{h}^{\pi_{k}}(s,\cdot),\pi_{s}\rangle|\leq\max_{s\in\mathcal{S}}\lVert\hat{Q}_{h}^{\pi_{k}}(s,\cdot)-Q_{h}^{\pi_{k}}(s,\cdot)\rVert_{\infty}\cdot\lVert\pi_{s}\rVert_{1}\leq b\,.$
@@
Note that the cardinality of each set $\mathcal{S}_{k}$ is evidently upper bounded by $|\mathcal{S}|$ but also by $|\mathcal{A}|^{h-k}\prod_{i\in\left[k+1,h\right]}M_{i}$ as the number of states in each set grows by a factor of at most $|\mathcal{A}|M_{k}$ at step $k$. In this section we prove some concentration bounds on the above estimators, which will be useful for the rest of the analysis.
@@
$|\hat{V}_{k+1}(\tilde{s})-({\mathcal{T}}\hat{V}_{k})(\tilde{s})|\leq C_{k+1}^{V}(\delta):=\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{A}|/\delta)}{M_{k}}}\,.$ (37)
@@
$|\hat{V}_{k+1}(\tilde{s})-({\mathcal{T}}\hat{V}_{k})(\tilde{s})|\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{A}|/\delta)}{M_{k}}}\,.$ (41)
@@
$\leq 2|\mathcal{A}|\exp(-2\frac{(1-\gamma)^{2}}{\gamma^{2}}t^{2}M_{k})\,.$ (42)
B.2.1 Estimation Error Bound for $\hat{Q}_{h}$
@@
$\max_{a\in\mathcal{A}}\lVert\hat{Q}_{k}(\cdot,a)-Q^{\pi}_{k}(\cdot,a)\rVert_{\mathcal{S}_{k}}\leq\frac{\gamma^{H+k}}{1-\gamma}+\frac{\gamma^{k}}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{S}_{0}|/\delta_{V})}{M_{0}}}+\frac{\gamma^{2}(1-\gamma^{k-1})}{(1-\gamma)^{2}}\sqrt{\frac{\log(2|\mathcal{A}||\mathcal{S}_{0}|(k-1)/\delta_{J})}{M}}$ (53)
@@
B.2.2 End of the Proof of Theorem 5.4
For any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, define the set $|\mathcal{S}_{h}|=\{s\}$ and the sets $|\mathcal{S}_{k}|$ accordingly for $10\leq k\leq h-$ 1. Here we set $M_{t}:=M>0\,,\forall t\in\left[1,h\right]$. Assuming the worst case growth of $|S_{k}|\leq|S_{k+1}||\mathcal{A}|M_{k+1}$, we can bound $|\mathcal{S}_{0}|$ deterministically as $|\mathcal{S}_{0}|\leq|\mathcal{A}|^{h}M_{h}M^{h-1}$. Using the bound in the previous section, it is possible to bound the error of $\hat{Q}_{h}(s,a)$ as follows:
@@
$\leq\frac{\gamma^{H+h}}{1-\gamma}+\frac{\gamma^{h}}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{S}_{0}|/\delta_{V})}{M_{0}}}+\frac{\gamma^{2}(1-\gamma^{h-1})}{(1-\gamma)^{2}}\sqrt{\frac{\log(2|\mathcal{A}||\mathcal{S}_{0}|h/\delta_{J})}{M}}\,,$ (54)
@@
$\forall 1\leq j\leq h\,,M_{j}:=M$ $>\frac{9\gamma^{4}(1-\gamma^{h-1})^{2}}{(1-\gamma)^{4}b^{2}}\log(2h|\mathcal{A}||\mathcal{S}_{0}||\mathcal{C}|/\delta_{J})\,,$ (55)
@@
Picking $\mathcal{C}=\mathcal{S}\times\mathcal{A}$ gives $\lVert\hat{Q}_{h}-Q^{\pi}_{h}\rVert_{\infty}\leq b$. If estimates for $\{V_{k}\}_{k=1}^{h}$ are reused across state action pairs, and the components of each value function are only computed at most once for every state $s\in|\mathcal{S}|$, then the overall number of samples is of the order of $KM_{0}H|\mathcal{S}_{0}|+Kh|\mathcal{S}_{0}|\cdot|\mathcal{C}|M$. In particular, with $\mathcal{C}=\mathcal{S}\times\mathcal{A}$, the number of samples becomes of the order of $KM_{0}H|\mathcal{S}|+Kh|\mathcal{S}|\cdot|\mathcal{A}|M$ where we reuse estimators (which are in this case computed for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$. Let $K>\frac{1}{h(1-\gamma)}\log(\frac{4}{\epsilon(1-\gamma)(1-\gamma^{h})})$ as in Theorem 5.4. Choosing $\delta_{V}=\delta_{J}=\delta/2$ for some $\delta\in(0,1)$ and $b=\frac{\epsilon(1-\gamma)(1-\gamma^{h})}{4}$ yields the following bound with probability at least $1-\delta$ (by Theorem 5.3):

### v2

@@ 8 Related Work
Policy Mirror Descent and Policy Gradient Methods. Motivated by their empirical success (Schulman et al., 2015, 2017), the analysis of PG methods has recently attracted a lot of attention (Agarwal et al., 2021; Bhandari and Russo, 2021; Khodadadian et al., 2022; Bhandari and Russo, 2024; Xiao, 2022; Lan, 2023; Johnson et al., 2023). Among these works, a line of research focused on the analysis of PMD (Xiao, 2022; Lan, 2023; Li and Lan, 2023; Li et al., 2023; Johnson et al., 2023; Alfano et al., 2023) as a flexible algorithmic framework, and its particular instances such as the celebrated natural policy gradient (Khodadadian et al., 2022; Yuan et al., 2023). In particular, for solving unregularized MDPs (which is our main focus), Xiao Xiao (2022) established linear convergence of PMD in the tabular setting with a rate depending on the instance dependent distribution mismatch coefficients, and extended their results to the inexact setting using a generative model. Similar results were established for the particular case of the natural policy gradient algorithm (Bhandari and Russo, 2021; Khodadadian et al., 2022). More recently, Johnson, Pike-Burke, and Rebeschini Johnson et al. (2023) showed a dimension-free rate for PMD with exact policy evaluation using adaptive stepsizes, closing the gap with PI. Notably, they further show that this rate is optimal and their analysis inspired from PI circumvents the use of the performance difference lemma which was prevalent in prior work. We improve over these results using lookahead and we extend our analysis beyond the tabular setting by using linear function approximation. To the best of our knowledge, the use of multi-step greedy policy improvement in PMD and its analysis are novel in the literature.
@@
Policy Iteration with Multiple-Step Policy Improvement. Multi-step greedy policies have been successfully used in applications such as the game of Go and have been studied in a body of works (Efroni et al., 2018a, b; Winnicki and Srikant, 2023; Efroni et al., 2020; Tomar et al., 2020). In particular, multi-step greedy policy improvement has been studied in conjunction with PI in Efroni et al. (2018a); Winnicki and Srikant (2023). Efroni et al. (Efroni et al., 2018a) introduced $h$-PI which incorporates $h$-lookahead to PI and generalized existing analyses on PI with 1-step greedy policy improvement to $h$-step greedy policies. However, their work requires access to an $h$-greedy policy oracle, and the $h$-PI algorithm is unstable in the stochastic setting even for $h=1$ because of potentially large policy updates. We address all these issues in the present work with our $h$-PMD algorithm. Winnicki and Srikant Winnicki and Srikant (2023) showed that a first-visit version of a PI scheme using a single sample path for policy evaluation converges to the optimal policy provided that the policy improvement step uses lookahead instead of a $1$-step greedy policy. Unlike our present work, the latter work assumes access to a lookahead greedy policy oracle. More recently, Alegre et al. Alegre et al. (2023) proposed to use lookahead for policy improvement in a transfer learning setting. Extending Generalized Policy Improvement (GPI) Barreto et al. (2020) which identifies a new policy that simultaneously improves over all previously learned policies each solving a specific task, they introduce $h$-GPI which is a multi-step extension of GPI. We rather consider a lookahead version of PMD including policy gradient methods instead of PI. We focus on solving a single task and provide policy optimization guarantees in terms of iteration and sample complexities.
@@
Lookahead-Based Policy Improvement with Tree Search. Lookahead policy improvement is usually implemented via tree search methods (Browne et al., 2012; Shah et al., 2020; Schrittwieser et al., 2019; Kocsis and Szepesvári, 2006b; Hallak et al., 2023; Efroni et al., 2019). Recently, Dalal et al. Dalal et al. (2022) proposed a method combining policy gradient and tree search to reduce the variance of stochastic policy gradients. The method relies on a softmax policy parametrization incorporating a tree expansion. Our general $h$-PMD framework which incorporates a different lookahead policy improvement step also brings together PG and tree search methods. Morimura et al. Morimura et al. (2022) designed a method with a mixture policy of PG and MCTS for non-Markov Decision Processes. Grill et al. Grill et al. (2020a) showed that AlphaZero as well as several other MCTS algorithms compute approximate solutions to a family of regularized policy optimization problems which bear similarities with our $h$-PMD algorithm update rule for $h=1$ (see sec. 3.2 therein).
@@ 9 Conclusion
In this work, we introduced a novel class of PMD algorithms with lookahead inspired by the success of lookahead policies in practice. We have shown an improved $\gamma^{h}$-linear convergence rate depending on the lookahead depth in the exact setting. We proposed a stochastic version of our algorithm enjoying an improved sample complexity. We further extended our results to scale to large state spaces via linear function approximation with a performance bound independent of the state space size. Our paper offers several interesting directions for future research. A possible avenue for future work is to investigate the use of more general function approximators such as neural networks to approximate the lookahead value functions and scale to large state-action spaces. Recent work for PMD along these lines (Alfano et al., 2023) might be a good starting point. Enhancing our algorithm with exploration mechanisms (see e.g. the recent work O’Donoghue (2023)) is an important future direction. Our PMD update rule might offer some advantage in this regard as was recently observed in the literature (Li and Lan, 2023). Constructing more efficient fully online estimators for the lookahead action values using MCTS and designing adaptive lookahead strategies to select the tree search horizon (Rosenberg et al., 2023) are promising avenues for future work. Our algorithm brings together two popular and successful families of RL algorithms: PG methods and tree search methods. Looking forward, we hope our work further stimulates research efforts to investigate practical and efficient methods enjoying the benefits of both classes of methods.
Acknowledgments and Disclosure of Funding
We would like to thank the anonymous area chair and reviewers for their useful comments which helped us to improve our paper. Most of this work was completed when A.B. was affiliated with ETH Zürich as a postdoctoral fellow. This work was partially supported by an ETH Foundations of Data Science (ETH-FDS) postdoctoral fellowship.
@@ Appendix A Additional Details and Proofs for Section 4: Exact $h$-PMD
$=({\mathcal{T}^{\pi_{k+1}}})^{n+2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\sum_{i=1}^{n}\gamma^{i+1}c_{k}\mathbf{e}+\gamma c_{k}\mathbf{e}$
$=({\mathcal{T}^{\pi_{k+1}}})^{n+2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\sum_{i=1}^{n+1}\gamma^{i}c_{k}\mathbf{e}\,,$ (12)
@@
A.3 Additional Comments About the Stepsizes in Theorem 4.1
About the cost of computing adaptive stepsizes. There is typically only one greedy policy at each iteration, and in the case that there is more than one it is possible to just pick any arbitrary greedy policy and the lower bound condition on the step size will still be satisfied. The cost of computing the step size is typically the cost of computing a Bregman divergence between two policies. We elaborate on this in the following using the discussion in Johnson et al. [2023] p. 7 in ‘Computing the step size’ paragraph which also applies to our setting. We use the step size condition $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in G_{h}(V^{\pi_{k}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$. We have a minimum in the condition which means we can just pick a single greedy policy $\tilde{\pi}\in G_{h}(V^{\pi_{k}})$ (which can easily be computed given a lookahead value function in the exact setting or its estimate in an inexact setting). As for the maximum over the state space, this is because we use the same step size in all the states. This condition can readily be replaced by a state dependent step size $\eta_{k}(s)\geq\frac{1}{c_{k}}D_{\phi}(\tilde{\pi}(\cdot|s),\pi_{k}(\cdot|s))$ which is enough for our results to hold.
About the condition on the stepsizes. First, notice that PMD has been analyzed using an increasing stepsize in the exact setting (see e.g. Lan [2023], Xiao [2022], Li et al. [2023]) and this is natural as this corresponds to emulating Policy Iteration. In the present work, notice that we do not require the stepsize to go to infinity and the stepsize does not typically go to infinity under our condition. Recall that we only require the stepsize to satisfy the condition: $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in\mathcal{G}_{h}(V^{\pi_{k}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}.$ As the policy gets closer to the optimal one it should also get closer to a greedy policy (the optimal policy is greedy with respect to its value function) and the term containing a Bregman divergence will vanish (since the optimal policy is already greedy with respect to its own value function). This effect balances the decreasing $c_{k}$. Notice that we typically choose $c_{k}=\gamma^{2h(k+1)}$.
We refer the reader to section D.4 for additional observations regarding stepsizes in experiments.
@@ Appendix B Additional Details and Proofs for Section 5: Inexact $h$-PMD
B.1 Details and Precise Notation for Lookahead Computation
We describe here our approach to estimate $Q_{h}^{\pi}$ in more detail and introduce the notation necessary for the analysis in the following sections. Assume that we are estimating $Q_{h}^{\pi}(s,a)$ at a given fixed state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$. We define a sequence $(\mathcal{S}_{k})$ of sets of visited states as follows: Set $\mathcal{S}_{h}=\{s\}$ and define for every $k\in[0,h-1]$, for every $\tilde{s}\in\mathcal{S}_{k+1}$ and for ever $\tilde{a}\in\mathcal{A}$ the multiset $\mathcal{S}_{k}(\tilde{s},\tilde{a}):=\{s_{j}^{\prime}\sim P(\cdot|\tilde{s},\tilde{a}):j\in[M_{k+1}]\}$ (including duplicated states) where $M_{k+1}$ is the number of states sampled from each state $\tilde{s}\in\mathcal{S}_{k+1}\,,$ and $\mathcal{S}_{k}=\bigcup_{\tilde{s}\in\mathcal{S},\tilde{a}\in\mathcal{A}}\mathcal{S}_{k}(\tilde{s},\tilde{a})\cup\mathcal{S}_{k+1}$ (without duplicated states). The estimation procedure is recursive. First we define the value function $\hat{V}_{1}$ by $\hat{V}_{1}(\tilde{s}):=\hat{V}^{\pi}(\tilde{s})\,,\,\forall\tilde{s}\in\mathcal{S}_{0}\,,$ where $\hat{V}^{\pi}(\tilde{s})$ is an estimate of $V^{\pi}(\tilde{s})$ computed using $M_{0}$ rollouts of the policy $\pi$ from state $\tilde{s}\,.$ More precisely, we sample $M_{0}$ trajectories $(\tau_{j}(\tilde{s}))_{j\in[M_{0}]}$ of length $H\geq 1$ rolling out policy $\pi$, i.e. $\tau_{j}(\tilde{s}):=\{(s_{k}^{j},a_{k}^{j},r(s_{k}^{j},a_{k}^{j}))\}_{0\leq k\leq H-1}$ with $s_{0}^{j}=\tilde{s}$ and $a_{k}^{j}\sim\pi(\cdot|s_{k}^{j})$. Then the Monte Carlo estimator is computed as $\hat{V}^{\pi}(\tilde{s})=\frac{1}{M_{0}}\sum_{j=1}^{M_{0}}\sum_{k=0}^{H-1}\gamma^{k}r(s_{k}^{j},a_{k}^{j})\,.$ For each stage $k\in\left[1,h\right]$ for $h\geq 1$, we define the following estimates for every $\tilde{s}\in\mathcal{S}_{k},\tilde{a}\in\mathcal{A}$,
$\hat{Q}_{k}(\tilde{s},\tilde{a})$ $:=r(\tilde{s},\tilde{a})+\frac{\gamma}{M_{k}}\sum_{s^{\prime}\in\mathcal{S}_{k-1}}\hat{V}_{k}(s^{\prime})\,,\quad\hat{V}_{k+1}(\tilde{s}):=\max_{\tilde{a}\in\mathcal{A}}\hat{Q}_{k}(\tilde{s},\tilde{a})\,,$ (22)
B.2 Proof of Theorem 5.3: Inexact $h$-PMD
@@ B.1 Details and Precise Notation for Lookahead Computation
$\lVert\hat{V}^{{\pi_{k}}}_{h}(\pi)-{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\rVert_{\infty}$ $=\max_{s\in\mathcal{S}}|\langle\hat{Q}_{h}^{\pi_{k}}(s,\cdot)-Q_{h}^{\pi_{k}}(s,\cdot),\pi_{s}\rangle|$
$\leq\max_{s\in\mathcal{S}}\lVert\hat{Q}_{h}^{\pi_{k}}(s,\cdot)-Q_{h}^{\pi_{k}}(s,\cdot)\rVert_{\infty}\cdot\lVert\pi_{s}\rVert_{1}\leq b\,.$
@@
Note that the cardinality of each set $\mathcal{S}_{k}$ is evidently upper bounded by $S$ but also by $A^{h-k}\prod_{i\in\left[k+1,h\right]}M_{i}$ as the number of states in each set grows by a factor of at most $AM_{k}$ at step $k$. In this section we prove some concentration bounds on the above estimators, which will be useful for the rest of the analysis.
@@
$|\hat{V}_{k+1}(\tilde{s})-({\mathcal{T}}\hat{V}_{k})(\tilde{s})|\leq C_{k+1}^{V}(\delta):=\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2A/\delta)}{M_{k}}}\,.$ (36)
@@
$|\hat{V}_{k+1}(\tilde{s})-({\mathcal{T}}\hat{V}_{k})(\tilde{s})|\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2A/\delta)}{M_{k}}}\,.$ (40)
@@
$\leq 2A\exp(-2\frac{(1-\gamma)^{2}}{\gamma^{2}}t^{2}M_{k})\,.$ (41)
B.3.1 Estimation Error Bound for $\hat{Q}_{h}$
@@
$\max_{a\in\mathcal{A}}\lVert\hat{Q}_{k}(\cdot,a)-Q^{\pi}_{k}(\cdot,a)\rVert_{\mathcal{S}_{k}}\leq\frac{\gamma^{H+k}}{1-\gamma}+\frac{\gamma^{k}}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{S}_{0}|/\delta_{V})}{M_{0}}}+\frac{\gamma^{2}(1-\gamma^{k-1})}{(1-\gamma)^{2}}\sqrt{\frac{\log(2A|\mathcal{S}_{0}|(k-1)/\delta_{J})}{M}}$ (52)
@@
B.3.2 End of the Proof of Theorem 5.4
For any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, define the set $|\mathcal{S}_{h}|=\{s\}$ and the sets $|\mathcal{S}_{k}|$ accordingly for $10\leq k\leq h-$ 1. Here we set $M_{t}:=M>0\,,\forall t\in\left[1,h\right]$. Assuming the worst case growth of $|S_{k}|\leq|S_{k+1}|AM_{k+1}$, we can bound $|\mathcal{S}_{0}|$ deterministically as $|\mathcal{S}_{0}|\leq A^{h}M^{h}$. Using the bound in the previous section, it is possible to bound the error of $\hat{Q}_{h}(s,a)$ as follows:
@@
$\leq\frac{\gamma^{H+h}}{1-\gamma}+\frac{\gamma^{h}}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{S}_{0}|/\delta_{V})}{M_{0}}}+\frac{\gamma^{2}(1-\gamma^{h-1})}{(1-\gamma)^{2}}\sqrt{\frac{\log(2A|\mathcal{S}_{0}|h/\delta_{J})}{M}}\,,$ (53)
@@
$\forall 1\leq j\leq h\,,M_{j}:=M$ $>\frac{9\gamma^{4}(1-\gamma^{h-1})^{2}}{(1-\gamma)^{4}b^{2}}\log(2hA|\mathcal{S}_{0}||\mathcal{C}|/\delta_{J})\,,$ (54)
@@
Picking $\mathcal{C}=\mathcal{S}\times\mathcal{A}$ gives $\lVert\hat{Q}_{h}-Q^{\pi}_{h}\rVert_{\infty}\leq b$. If estimates for $\{V_{k}\}_{k=1}^{h}$ are reused across state action pairs, and the components of each value function are only computed at most once for every state $s\in S$, then the overall number of samples is of the order of $KM_{0}H|\mathcal{S}_{0}|+Kh|\mathcal{S}_{0}|\cdot|\mathcal{C}|M$. In particular, with $\mathcal{C}=\mathcal{S}\times\mathcal{A}$, the number of samples becomes of the order of $KM_{0}HS+KhS\cdot AM$ where we reuse estimators (which are in this case computed for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$. Let $K>\frac{1}{h(1-\gamma)}\log(\frac{4}{\epsilon(1-\gamma)(1-\gamma^{h})})$ as in Theorem 5.4. Choosing $\delta_{V}=\delta_{J}=\delta/2$ for some $\delta\in(0,1)$ and $b=\frac{\epsilon(1-\gamma)(1-\gamma^{h})}{4}$ yields the following bound with probability at least $1-\delta$ (by Theorem 5.3):